OpenStudy (e.cociuba):

Solve the equation using the quadratic formula. 2m^2 + 2m - 12 = 0

OpenStudy (campbell_st):

the formula is \[m = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you have a = 2, b = 2 and c = -12 just substitute and evaluate.

OpenStudy (e.cociuba):

Ok I got it. Thanks! :)